### v1

# Emergent interactions lead to collective frustration in robotic matter

Abstract
Current artificial intelligence systems show near-human-level capabilities when deployed in isolation. Systems of a few collaborating intelligent agents are being engineered to perform tasks collectively. This raises the question of whether robotic matter, where many learning and intelligent agents interact, shows emergence of collective behaviour. And if so, which kind of phenomena would such systems exhibit? Here, we study a paradigmatic model for robotic matter: a stochastic many-particle system in which each particle is endowed with a deep neural network that predicts its transitions based on the particles’ environments. For a one-dimensional model, we show that robotic matter exhibits complex emergent phenomena, including transitions between long-lived learning regimes, the emergence of particle species, and frustration. We also find a density-dependent phase transition with signatures of criticality. Using active matter theory, we show that this phase transition is a consequence of self-organisation mediated by emergent inter-particle interactions. Our simple model captures key features of more complex forms of robotic systems.
The recent advances in artificial intelligence have brought about systems with near-human-level capabilities in some specific tasks. These advances mostly rely on deep neural networks, which can be trained efficiently using large data sets. In robotics, artificial intelligence is used to give robots the capability to interpret and manipulate their environment [1]. Recently, artificial intelligence systems are being combined to achieve better performance compared to individual systems [2, 3, 4, 5, 6]. In the context of robotics, such collectives contain multiple robotic units that are programmable and interact with their environment. Programmed interactions enable such robot collectives to create different formations [7, 8, 9] and react to global signals to manipulate objects [10].
Figure 1: Figure 1. (a) Feedback between collective particle dynamics and the evolution of deep neural networks in stochastic many-particle systems driven by deep learning. (b) One-dimensional lattice gas implementation. (c) Numerical protocol of stochastic updating and learning.
Robot collectives have been experimentally implemented across spatial scales and system sizes  [11, 12]. Due to the limitations imposed by miniaturization, there is a trade-off between the number of robots in a collective and the capacity of artificial intelligence that can be implemented in individual units. Macroscopic robot collectives comprise, for example, small groups of interacting robots which are each controlled by a performant artificial intelligence [13]. Smaller-scale robots are controlled by microcontrollers and can interact with thousands of robots for self-assembly, collective transport, collective perception, and more [7]. At the micron scale, robots interact via physical or chemical interactions between particles that are either self-driven active particles or particles driven by external forces and controlled via global signals. Such systems have, for example, been implemented as magnetic micro-disks driven by time-varying magnetic fields. Due to their small scale, the capabilities of individual particles in these systems are limited and they are therefore usually controlled externally using global signals.
Conceptually, one can distinguish between different levels on which learning and execution occurs. In decentralised learning and decentralised execution, units learn and behave independently of each other [14]. In the case of centralised learning and decentralised execution, units are controlled by a single machine-learning algorithm, which has been employed in active matter systems to learn chemotactic behaviour [15, 16]. Finally, in the case of centralised learning and decentralised execution, where a single neural network is used by all particles individually but trained using the trajectory of all particles, 2D and 3D active intelligent swimmers moving collectively to reduce their energy expenditure  [17]. Empirical studies in robot collectives have highlighted the emergence of subgroups of robots that may fulfil specialized tasks [18].Such speciation phenomena include the emergence of implicit leaders that provide direction of guidance to other robots. Such experiments have also highlight the emergence of deadlock states, robots perform suboptimal behaviour [19, 20, 21].
Here, we study a paradigmatic theoretical framework for robotic matter that predicts key phenomena of empirical robot collectives. Specifically, we study stochastic many-particle systems in which each particle individually is endowed with artificial intelligence. In these systems the stochastic state transitions are defined by deep neural networks associated with each particle. The predictions by distinct deep neural networks are based on a particle’s environment and they learn by analysing the particle’s trajectory over time. We find that these systems exhibit complex spatio-temporal behaviour: Over time, increasingly complex interactions between particles evolve in a succession of temporal learning regimes. We observe the evolution of multiple particle species and frustration, where the system is locked long-term in a suboptimal state. We also find a density-dependent phase transition with signatures of criticality. Using active matter theory, we show that this phase transition exhibits signatures of criticality and is a consequence of self-organisation mediated by emergent interactions between particles.
Figure 2: Representative trajectories showing the positions of particles over time in distinct temporal regimes (rows). Each column corresponds to a simulation at the specified particle density. For visual purposes we rescaled positions and time in the following way: The x-axis is the lattice position divided by the total lattice size (from left to right 2e4, 2e3, 1e3, 5e2, 4e2, 334 sites, respectively). The y-axis denotes the time in episodes divided by the typical time it takes for a particle to cross the system once. The starting time for each row is indicated in grey and the final time is on the top-right corner of each panel. Line colors show the sign of the average relative velocity of that particle in the last 2e5 episodes of the simulations. Trajectories of representative particles are highlighted in bold. For densities 0.4, 0.5, and 0.6 we have randomly sampled 100 particles to make individual lines distinguishable.
Results
Model definition
To begin, we consider a stochastic many-particle system where each particle, indexed by $i$, is described by a random variable, $x_{i}$, Fig. 1(a). The state of the system at a given time is then defined by the states of all $N$ particles, $\{x_{i}\}_{i=1,\ldots,N}$. The states of particles are updated according to transition rates between two states $\{x_{i}\}_{i=1,\ldots,N}$ and $\{x_{i}^{\prime}\}_{i=1,\ldots,N}$ per unit time. These transition rates follow from the probabilities of each particle performing individual state transitions $x_{i}\to x_{i}^{\prime}$. In our model, each particle is endowed with a unique deep neural network (DNN) which takes the system configuration, $\{x_{i}\}_{i=1,\ldots,N}$, as an input and computes the single-particle transition rate as an output. Because the system configuration includes the states of other particles, the DNNs may encode effective interactions between particles. DNNs can, in theory, approximate any continuous functions up to any desired accuracy [22], such that the single-particle transition rates may, in principle, encode arbitrarily complex interactions. These interactions may change over time as particles learn from observing their past trajectories and update the parameters of their DNNs based on a reward function. Taken together, the particles in our model are endowed with sensors, processors and actuators  [11]. The model describes a multi-scale system, in which DNNs determine the dynamics of individual particles and, vice versa, collective states on the macroscopic scale influence the learning of the parameters of DNNs. This gives two-way feedback between collective states emerging from effective interactions between particles on the macroscopic scale and the dynamics of DNN parameters,Fig. 1(a).
Here, to investigate the range of possible phenomenologies such systems may exhibit, we study the arguably simplest possible implementation of this framework. In this implementation, the state $x_{i}$ of particles is defined by their position on a one-dimensional lattice of $L$ sites with periodic boundary conditions, Fig. 1(b). A given particle $i$ can hop to the right neighbouring site with probability $p_{i}$ and to the left neighbouring site with probability $1-p_{i}$. This fixes the unit of time. These transition rates are determined by the output layer of feed-forward DNNs which take the lattice environment of size $l$ centered around a given particle as an input. The DNN parameters associated with a given particle are dynamically updated using deep reinforcement learning on the particles’ past trajectory to maximise a discounted reward. The reward function gives a large negative reward when a particle moves to an already-occupied lattice site, and a small positive reward otherwise (Appendix  A).
We used the kinetic Monte Carlo method to perform stochastic simulations of this model  [23]. To reduce the computational cost of training the DNNs, we divided the simulation time into intervals, termed episodes, Fig. 1(c). During each episode, we updated the positions of particles while keeping the DNN parameters fixed. At the end of each episode, we updated the values of the DNN parameters using reinforcement learning. There are different algorithms for training a collection of agents: Multi-agent reinforcement learning (MARL) is used when multiple agents are in a common environment and should learn to cooperate or compete with each other  [24, 25, 26, 27, 28]. Group-agent reinforcement learning (GARL) is used for cases in which each agent operates in a separate environment and only communicates with others to share knowledge [29, 30]. Here, we used an actor-critic reinforcement learning scheme with temporal-difference learning scheme to update the DNN parameters by maximising the cumulative discounted reward calculated from the particle trajectories in each episode  [31] (see Methods). All figures shown here were obtained from simulations, where the number of particles, $N$, was set to 200200200, the local environment size, $l$, was 212121, the system size, $L$, was set depending on particle density, and the discount factor, $\gamma$, was 0.90.90.9. DNNs had a triangular architecture with 21 neurons in the input layers, and 10, 5, and 2 neurons in the successive layers, respectively. We simulated up to $5\cdot 10^{6}$ episodes with learning rate $5\cdot 10^{-4}$. We also tested DNNs with up to five layers, but we observed no qualitative differences in the simulations. To simulate dilute systems without reducing the number of particles, we increased the number of lattice sites to achieve a given particle density while keeping the size of the environment, $l$, fixed. This was to ensure that the DNN sizes and the training dynamics at different particle densities are consistent. This fixes the units of length to the lattice spacing.
From the definition of the model, several questions must be asked that in this form do not arise in active matter systems: First, particles appear in other particles’ environment. Does this lead to the emergence of effective interactions between particles? If so, then, secondly, deep neural networks can, in principle, represent arbitrarily complex interactions that might be qualitatively different for each environment and each particle. Are robotic matter systems amenable to physics-type descriptions? Finally, would effective interactions between particles lead to collective behaviour as in other non-equilibrium systems studied in physics?
Dynamical regimes and density-dependent behaviour
Figure 2(a) shows representative trajectories of simulation results for different particle densities and different time windows in a given simulation. Despite the simplicity of the model, we find that particles form complex spatio-temporal structures after a short transient time. These structures change qualitatively over time, indicating that the dynamics transits through different temporal regimes. The spatio-temporal structures also depend on the particle density in that particles at low densities tend to align their velocity in the same direction, while above a density threshold of 0.3 particles at high densities move in opposite directions. In the following, we will analyse these changes as a function of the density more rigorously.
Figure 3: (a) Average velocity of particles relative to the average velocity of all particles in a simulation. Particles are grouped by their asymptotic velocity, which is calculated as an average over the final 2e5 episodes of the simulations. The dashed lines on the time axis denote in this and in all following figure two characteristic times: $t_{1}\approx 1e4$ episodes marks the initial formation of global alignment at low densities; $t_{2}\approx 3e4$ episodes marks the initial formation of two groups of particles moving in opposite direction at high densities. (b) Average absolute velocity of all particles for two representative densities (colours defined in (c)). (c) Log2-fold change of the standard deviation of the number of right-moving particles across simulations. The lower horizontal dashed line is the variance expected if the direction of movement was assigned by chance. The upper dashed line is the maximal variance expected for globally aligned motion. (d) Average reward for two representative densities as a function of time. The inlay is a magnification that shows a decrease in the reward for density 0.5. Colours as in (c).
To this end, we first consider the relative velocity of particles with respect to each other. It is defined by the difference between the individual, instantaneous velocity of a given particle, $v_{i}$, and the average velocity in the system at a given time, $\expectationvalue{v_{i}}$, $v_{i}-\expectationvalue{v_{i}}$. We calculated the relative velocity for each particle and averaged over particles with the same sign of the asymptotic velocity. Figure 3(a) shows that, for low densities, the particles have vanishing relative velocities and therefore are moving in the same direction with a similar absolute velocity. Figure 3(b) shows that this occurs at a characteristic time $t_{1}\approx 1e4$ episods. For densities larger than 0.3, particles form two groups moving in opposite directions. This could reflect a random assignment of the direction of movement of particles. However, the variance of particles moving in a given direction across all simulations is smaller than what is expected from a statistically independent assignment and predicted by the binomial for millions of episodes ($p=0.003$, F-test, Fig. 3(c)). Significantly, this implies that, in the high-density regime, the direction of particles is not chosen randomly, but tuned into a precisely balanced proportion of left- and right-moving particles on the system level.
The fine-tuning of particle directions to a 50-50 split at high densities seems counterintuitive given that they are rewarded for not colliding with other particles. Indeed, while the average reward at low concentrations increases monotonically and reaches an optimal value, at high densities the reward first increases and then decreases at a time that is comparable to the formation of the two groups of particles, Fig. 3(d).
This observation raises two questions: first, what is the reason for the abrupt change in the behaviour of particles? And secondly, why is the high-density regime fine-tuned to a suboptimal reward? A potential answer to these questions could either be that particles learn different behaviours at high and low densities. This should then be reflected in the dynamics of the parameters of the DNNs. An alternative explanation could be that the abrupt change in behaviour emerges from self-organisation due to interactions between particles. In the following, we will aim to understand the origin of this behaviour by investigating in detail the spatio-temporal structures, the time evolution of neural networks, and emergent interactions between particles.
Spatio-temporal structures
Figure 4: (a) Average distance to the nearest-neighbour particle rescaled by the nearest-neighbour distance (NND) expected for a spatially uniform distribution of particles (horizontal dashed line). Vertical dashed lines denote characteristic times defined in Fig. 3. (b) Fraction of particles sharing a lattice site with at least one other particle rescaled by the expected value for a uniform distribution (horizontal dashed line). (c) Heatmaps of the pair correlation function for low and high particle densities. The pair correlation function is rescaled by the average density of particles (grey).
To quantify the time evolution of spatial correlations between particle positions, we calculated the average distance to the nearest-neighbour particle over time, Fig. 4(a). In both density regimes, the nearest-neighbour distance decreases initially with time, showing that particles tend to cluster with other particles. The time evolution of the average nearest-neighbour distance is highly non-monotonic, reflecting different temporal regimes. Despite the formation of clusters, particles nevertheless reduce the probability of overlapping with other particles, Fig. 4(b). To further understand the time evolution of spatial structures we calculated the pair correlation function, which gives the average density of particles as a function of the distance to a given particle, Fig. 4(c). It shows that particles indeed cluster, but the sites adjacent to a given particle tend to be unoccupied. The pair-correlation function also shows that particle-particle correlations are longer in range in the low-density regime compared to the high-density regime.
Time evolution of neural network parameters
To test whether the formation of spatio-temporal structures and the change of particle behaviour with the density is correlated to patterns in the time evolution of the DNNs, we calculated probability density functions of the parameters defining the output of DNNs, namely the weights and biases, over time. Initialised with a Gaussian distribution, the distributions of DNN parameters develop different degrees of multi-modality over time, reflecting that DNN parameters converge to defined values, Fig. 5(a). The modality of the distributions of DNN parameters changes at the characteristic times defining changes in the spatio-temporal behaviour, cf. Figs. 3 and 4. This suggests the existence of different learning regimes, where particles learn different input-output relations.
Figure 5: Probability density functions of neural network parameters (black lines) for logarithmically spaced time points. Density functions are shown separately for two different densities, the actor and critic networks constituting the deep reinforcement learning algorithm and the weights and biases in the neural networks.
Analysis of interactions
To understand how the dynamics of the DNN parameters relate to the formation of spatio-temporal structures, we analyse the forces the DNNs encode. Intuitively, interactions between particles are encoded in the temporal change in its velocity, averaged over all environments. To develop this idea, we first represent a given environment by a vector $\vec{e}$, which contains the number of particles at each position in the environment. We now quantify the change in the velocity due to changes in the particle’s environment. To this end, we note that the change in the velocity can be written as the product of changes due to changing environments, and the temporal evolution of the distribution of finding a given environment. Interactions are encoded in the former part, which is given by the directional gradient of the velocity with respect to changes in the environment. In our case, the velocity is determined by the probability of hopping right in a given environment via $2p(\vec{e})-1$. Therefore, the directional derivative of the DNN output for a given input is a natural choice for quantifying interactions. To implement this, for a given environment $\vec{e}_{0}$, we quantified the change in the transition rates $p_{i}$ when the environment $\vec{e}_{0}$ is changed slightly by $\Delta e$. Here, we perturb the environment $\Delta\vec{e}$ by the addition of a single particle. With this, we can quantify interactions by calculating the difference in the neural network output between the original environment and the perturbed environment,
$\frac{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}{\left|\Delta\vec{e}\right|}\,.$ (1)
To quantify the type of interactions more qualitatively, we multiply (1) with the sign of the distance of the particle in $\Delta e$ to obtain a quantity that quantifies the degree with which particles are repulsed by the addition of a new particle to their environment (repulsiveness), $R=-\operatorname{sign}{(\Delta x)}\left[{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}\right]\big{/}{\left|\Delta\vec{e}\right|}$. Negative values of the repulsiveness indicate attractive interactions.
Figure 6: (a) Schematic showing the computation of interactions. (b) Dimensionally reduced representations of interactions (UMAPs) for two representative densities (top and bottom). Each point represents one particle at a given simulation time and simulation run. Points that are close in the UMAP representation behave similarly in interactions. Colours signify simulation time, (c) simulation index, and (d) asymptotic velocity calculated in the final 2e5 episodes of the simulations.
To quantify the so-defined interactions for each particle, we sampled 1000 random environments from the set of environments that host at most 6 particles. We passed these environments, and the corresponding environments that have one additional particle, to the DNNs of each particle at multiple time points, record the DNNs’ output, and calculate the value of $R$, Fig. 6(a). Applying this procedure to all 1000 randomly-chosen environments, particles, and time points over 50 replicates per density leads to data with roughly $10^{7}$dimensions. As a first step towards quantifying interactions between particles, we therefore used a nonlinear dimensionality-reduction technique, uniform manifold approximation, and projection (UMAP), to obtain a first insight into the time evolution of interactions. Particles that are close in these projections tend to behave similarly in interactions. The UMAPs show the systematic time evolution of the response of particles to changes in their environments, Fig. 6(b). This time evolution is qualitatively different in the high and low density regimes, meaning that particles develop different interactions in both regimes, Fig. 6(b). All structures in the UMAPs are independent of the simulation run, Fig. 6(c), showing that particles consistently develop similar possible interactions across runs.
At high densities, we observe the emergence of four clusters of particles in the projected space after a time $2\cdot 10^{5}$ which corresponds to the formation of the two groups of particles moving in different directions, Fig. 2(c). This means that particles evolve into four distinct species, which are distinguished by the way they interact with their environments. Two pairs of these clusters identify particles moving ballistically in a given direction, Fig. 6(c). Particles moving in the same direction further separate into two groups, which interact in different ways depending on higher-order properties of the environments. We were not able to identify simple rules that distinguish between interactions of particles in the clusters corresponding to particles moving in the same direction, showing that particles develop complex interactions that do not follow patterns as in other systems studied in statistical physics.
To gain deeper insights into these interactions, we parametrised environments by the distance of the particle in $e_{j}$ , $\Delta x$, and the distance of the first and second nearest neighbours in $\vec{e}_{0}$, $k_{1}$ and $k_{2}$, respectively. Figure 7(a) shows that interactions are predominantly repulsive with a tendency for their strength to increase over time. This is expected by the choice of the reward function. Additionally, interactions tend be become longer in range over time, as particles obtain non-zero repulsion with respect to other particles at increasing distances. While particles at high densities only interact with the nearest and next-nearest neighbour sites particles at low densities interact over longer distances at large times. Figure 7(a) also shows characteristic branching points, that mark time points after which the response of particles in a given set of environments becomes non-homogeneous. At these time points, particles learn to distinguish an increasing number of different environments.
To investigate how interactions depend on the base environment $\vec{e}_{0}$ we further parametrised the base environment by the distance to the first nearest neighbour, $k_{1}$, and the distance to the second nearest neighbor, $k_{2}$. At low densities, particles first distinguish between base environments, where $k_{1}=1$ and $k_{2}$ is either equal or larger than 2, Figure 7(b). For low densities, after the emergence of nearest-neighbour repulsion, particles first learn to respond to the presence of a particle at a neighbouring site ($\Delta x=1$) depending on the position of the second nearest-neighbour particle. Specifically, particles learn to distinguish between environments, where the second nearest neighbour is two sites away compared to a longer distance. Later, particles learn to respond to the presence of a particle two sites away ($\Delta x=2$) depending on the position of a second nearest neighbour at the same distance. For high densities, after the emergence of nearest-neighbour repulsion, we observe a similar sequence of branching points but at later time points during the simulation.
A large number of environments tend to follow similar trends, indicating that the DNNs in particles do not distinguish between these environments in terms of repulsiveness and that DNNs initially do not interpret the full complexity of the environment. At characteristic times that partially overlap with the boundaries of the temporal regimes defined on the macroscopic scale we observe branching points after which environments become distinguishable by the DNNs to increasing degrees. Therefore, over time, DNNs of particles learn to distinguish between an increasing number of environments at characteristic time points. In between these time points, the form of interactions tends to be stable over increasingly long time periods.
Figure 7: To understand how interactions involve over time, we computed the time evolution of the repulsiveness, averaged over all particles. Each line shows the repulsiveness in a different environment. (a) Repulsiveness colored by the distance to the particle added to the base environments. (b) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour. (c) Repulsiveness for the case that a particle is added at distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour.
Phenomenological theory
With this, we are now in a position to come back to the question of what is the origin of the abrupt change in the behaviour of particle as a function of the particle density and the associated decrease of rewards. Our analyses have shown both: macroscopic order in space and time, which indicates collective behaviour of particles; and highly complex, density-dependent interactions that depend on evolving particle “species”. To test the hypothesis that the abrupt change in the behaviour of particles is a consequence of self-organization, we now develop a minimal, phenomenological model that only comprises a minimal set of interactions without explicit density dependence.
To this end, we use the framework of active matter theory, and Langevin equations. We consider the position $x_{i}(t)$, and velocity $v_{i}(t)$ of a particle $i$ at a given instance of time, $t$. The time evolution of the position and velocity of this particle is then described by a set of Langevin equations of the form
$\dot{x}_{i}=$ $v_{i}\,,$ (2)
$\dot{v}_{i}=$ $-\gamma(v_{i})v_{i}+K_{i}+\sqrt{2D}\eta_{i}\,.$ (3)
where $\gamma(v)$ is a generalized, non-equilibrium friction, $K_{i}$ is the force acting between particle $i$ and all other particles in its environment, $D$ is the noise amplitude, and $\eta$ is Gaussian white noise.
We first identify the shape of the interactions $K_{i}$ from our simulations. To this end, we take a coarse-grained perspective and quantify the force acting on a particle $i$ for a given average velocity field $u(x)$ and density field $\rho(x)$ averaged in a small environment around the position $x_{i}$, $K_{i}(\rho,u)$. In our simulations, interactions evolve over time and are highly complex. They cannot be easily translated into simple interaction rules. For the purpose of our argument, we identify two density-, particle, and species-independent components of interactions: First, we have shown in Fig. 7(b) that interactions between particles are predominantly repulsive. This gives rise to a contribution to the interaction term of the form $-\partialderivative*{\rho(x)}{x}\evaluated{}_{x=x_{i}}$.
Further, interactions aim to minimise not only the rate of collisions, but also their correlations. This is because particles maximise a discounted reward of the form $G=\sum_{j}\gamma^{j}r_{j}$ that depends not only on the instantaneous rate of collisions, but also on how they are correlated over time. Specifically, the expected discounted reward, $\langle G\rangle$ increases if the correlations in collisions decrease (Appendix D),
$\langle G\rangle\sim-\tau_{-}\langle\rho\rangle\left(1-\langle\rho\rangle\right)\left(1-r_{-}\right)+\mathcal{O}(\tau_{-}^{2}),$ (4)
where $\tau_{-}$ and $\tau_{+}$ are the autocorrelation times of collisions and non-collisions, respectively. The minimisation of the autocorrelation time is empirically supported by our simulations (Fig. S1). The autocorrelation time is proportional to the time a pair of particles spend in each other’s vicinity, such that particles minimise a function that is proportional to $1/|v-u|$. This then leads to a force that is given by the derivative of this term with respect to $v$, $\operatorname{sign}(v-u)/|v-u|^{2}$. Because the correlation scales with the density, the second contribution to the interaction terms takes the form $\rho\operatorname{sign}(v-u)/|v-u|^{2}$. Taken together, in our phenomenological description, this gives an effective force term for particle $i$ of the form
$K_{i}=\zeta\rho\frac{\operatorname{sign}(v-u)}{|v-u|^{2}}-\delta\partialderivative{\rho}{x}\evaluated{}_{x=x_{i}}\,.$ (5)
Figure 8: a) Representative trajectories showing the positions of particles at the specified particle density in the last 200 episodes. The colors show the sign of average velocity of the particles in the last 1e3 time points: red for positive and blue for a negative average velocity. b) Phase portrait of Eqs. (9) and (10) for density 0.1. c) phase portrait of Eqs. (9) and (10) for density 0.5. The colors show the sign of the velocity of the test particle in the Hartree approximation. d) velocity autocorrelation time obtained from simulations of Eqs. (7) and (8). The dashed line shows the density 0.3. e) Velocity autocorrelation time obtained from simulation of the full model in Fig. 1(b).
We now define the active friction term $\gamma(v)$. Because our system is out of equilibrium, the fluctuation-dissipation theorem does not hold, such that friction is not coupled to the noise amplitude. Instead, this term comprises a part that describes directed motion and a part that describes friction. A phenomenological term that describes both is the Rayleigh-Helmholtz friction [32], which has been used to describe a variety of active matter systems. It is of the form
$\gamma(v)=-\alpha v+\beta v^{3}\,.$ (6)
Taken together, in non-dimensional form, the time-evolution of velocity and position of a given particle reads
$\dot{x}=$ $v,,$ (7)
$\dot{v}=$ $v-v^{3}+s\frac{\rho\operatorname{sign}{(v-u)}}{(v-u)^{2}}-r\partialderivative{\rho}{x}+\sqrt{B}\eta,$ (8)
where we have dropped the index $i$. We have non-dimensionalised the Langevin equations by rescaling time and space by $t^{\prime}=\alpha t$ and $x^{\prime}=x\sqrt{\alpha\beta}$, respectively, and removed the apostrophes to simplify the notation. The dimensionless parameters $s=\zeta\sqrt{\beta^{5}/\alpha^{3}}$ and $r=\delta\sqrt{\beta^{3}/\alpha}$ quantify the strength of interactions due to the minimisation of correlations and repulsion relative to the effect of self-propulsion, respectively. $B=D\beta/\alpha^{2}$ is the dimensionless noise strength.
In this model, the relative strength of the two interaction terms implicitly depends on the particle density. For high densities, fluctuations in the particle density are weak, such that we expect the term that minimises correlations to dominate. Intuitively, this leads to a maximisation of the difference between the particle’s velocity $v$ and the average velocity $u$, which could globally lead to a split as observed in our simulations. For low densities, the term that minimises correlations is negligible and we expect particle repulsion to be dominant.
To test this numerically, we performed stochastic simulations of Eqs. (7) and (8) using a finite difference method (see Appendix C). We indeed found that particles show global alignment at low densities and a split into groups moving in opposite direction for high densities (Fig. 8). Again, the variance of particles moving in the positive direction across simulations is smaller than expected by chance ($p=2.22e^{-16}$), suggesting a collective tuning of the sizes of both groups. These observations suggest that the abrupt change in the behavior of particles in Fig. 3(a) is a phase transition resulting from collective behavior induced by interactions between particles.
To investigate this more quantitatively, we analytically derived the coevolution of the velocity of a particle, $v$, and the average velocity of all particles $\bar{u}$. To this end, followed Ref. [33] and used a moment expansion up to second order in the average velocity to obtain a set of nonlinear differential equations,
$\dot{v}=$ $(1-v^{2})v+s\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+r(\bar{u}-v),,$ (9)
$\dot{\bar{u}}=$ $\bar{u}-\bar{u}(\bar{u}^{2}+3T).$ (10)
Here, $T=\langle(v-\bar{u})^{2}\rangle$ is the variance of the velocities, which we estimated from our simulations of the Langevin equations. The co-evolution of the velocity of individual particles and the average velocity is summarized in phase portraits (Fig. 8). The phase portraits show attractors that correspond to asymptotic dynamical regimes of Eqs. (7) and (8). For low densities, we find two attractors with $v=\bar{u}$ that correspond to states, where the velocity of particles aligns with the average velocity in the system. For high densities, these attractors are located horizontally on a line given by $\bar{u}=0$. Therefore, the phenomenological model predicts the evolution of individual velocities to take specific positive or negative values under the constraint of a vanishing global velocity. Because the line $\bar{u}=0$ is stable, fluctuations around it are suppressed by feedback mechanisms. This corresponds to the observed state with precisely tuned equal proportions of particles moving in either direction.
Taken together, our analysis shows that the abrupt change of the behaviour as a function of the particle density is is an emergent phenomenon resulting from interactions between particles. It is therefore a phase transition as in statistical physics systems. For high densities, interactions between particles have the macroscopic effect . In this regime, particles optimise their reward locally, without being able to predict the suboptimal consequences of their behaviour on the macroscopic scale. This resembles a learning analogon to frustration phenomena in disordered systems or the fallacy of the commons in social sciences. This analysis raises the question of whether this phase transition resembles known phase transitions in (non-equilibrium) statistical physics. To investigate this, we calculated autocorrelation functions of the velocities of individual particles using the batch means method [34], both in the Langevin simulations and in the simulations of the full model comprising intelligent agents, for different particle densities. In both cases, the autocorrelation time diverges at the density at which the phase transition occurs. Diverging correlations are characteristic of critical points, suggesting critical behaviour. A further characterisation of the phase transition requires the computation of critical exponents, which is currently not feasible for intelligent agents. It is, however, of note that the phase transition occurs at a density, at which the range covered by the repulsive interactions of each particle multiplied by number of particles equals the lattice size, that is, the effective volume fraction is equal to 1. This could indicate a relation to jamming transitions.
Discussion
Taken together, we studied a paradigmatic model of robotic matter. In this model, particles in a stochastic lattice gas are endowed with deep neural networks that take the spatial environment of a particle as an input and predict transition rates. Our extensive simulations reveal a rich phenomenology, including complex emergent phenomena leading to a phase transition, the evolution of multiple particle species, temporal learning regimes, and a phenomenon akin to frustration, where particles adopt suboptimal strategies. The phase transition we observe shows signs of criticality. Although it is not possible to determine critical exponents, the behaviour close to the phase transition point shows similarities to a jamming transition.
Stochastic many-particle systems as the one studied in this work differ qualitatively from systems typically studied in active matter theory: because each particle has many more intrinsic degrees of freedom than the number of particles one can typically simulate on computers, they effectively lack microscopic symmetries. While in statistical physics one is typically interested in the breaking of microscopic symmetries, here the question arises if symmetries emerge as a consequence of learning, and whether these symmetries are then broken again. Because deep neural networks of sufficient size can, in principle, approximate arbitrarily complex function interactions between particles do not necessarily follow simple patterns. They may be qualitatively different between particles and environments. This raises the question of in how far physical theories are adequate in describing the time evolution of such systems. In our work, particles indeed developed complex interactions that cannot be described by simple mathematical functions. However, a simple phenomenological theory focussing on rudimentary aspects of these interactions is able to predict a key aspect of the system, namely the existence of a density-dependent phase transition.
Although our model is a drastic simplification of real robotic systems, our results show similarities to much more complex robotic systems. For example, Ref. [19] studied a hide and seek game between intelligent agents, in which agents passed through different, metastable learning regimes and interactions between agents lead to a decrease in the collected reward. Apart from interacting robots, also virtual artificial intelligence systems, like the widely-used large-language systems, are now beginning to be combined in order to achieve better performance [35]. This raises the question of how to design these systems to achieve an expected outcome. In some cases, interacting learning objectives straightforwardly translate into collective behaviour [14]. Our work implies that learning can be drastically influenced in unpredictable manners by collective behaviour on the macroscopic scale.
Acknowledgements
We thank Alexander Ziepke, Stefano Bo, Frank Jülicher, and the entire Rulands group for helpful discussions.
Author contributions
A.A. and O.B. contributed equally. A.A., O.B., and S.R. conceptualized the work. A.A. and S.R. supervised the work; O.B. wrote the simulation code and analysed O.B. performed analytical calculations. S.R. acquired funding. O.B. and S.R. wrote the initial draft of the manuscript. All authors edited and approved of the final version of the manuscript.
A Implementation of the stochastic simulations
[…]
$\sum_{i=1}^{n_{j}-1}-A_{i}\log{p_{i}},$
where $n_{j}$ is the number of transitions the particle $j$ made during the episode, and $A_{i}:=R_{i}+\gamma V_{i+1}-V_{i}$, and $p_{i},R_{i},V_{i}$ are the probability with which the particle made the transition, the reward it got and the value predicted by the critic at step $i$, respectively. To train the critic, we used the loss $\sum_{i=1}^{n_{j}-1}A_{i}^{2}/(n_{j}-1)$.
Figure S1: The average autocorrelation time of collisions over 101010 replicates. The shaded area around the curves denote the standard error. The averages are calculated at logarithmically samples time points.
B Numerical solution of the Langevin equations
We used the Euler scheme to integrate the Langevin equations. We used 200200200 particles with time increment $\delta t=0.01$, and the system size was determined based on $L=200/\langle\rho\rangle$ where $\langle\rho\rangle$ is the average particle density. We used periodic boundary conditions. We set $s=1.17,r=12.50,B=0.02$, and we used a local radius of 555 to calculate the local density field $\rho(x)$ and the local velocity field $u(x)$ for each particle. We run 151515 to 242424 replicates per particle density. To calculate the gradients of the local fields, we subtracted value of the field in the right of the particle at distance 555 from the value of the field in the left of the particle at distance 555, and divided this by 555.
C Non-dimensionalisation of the Langevin equation
To non-dimensionalize the Langevin equations, we rescaled time and positions as
$t^{\prime}$ $:=\alpha t$
$x^{\prime}$ $:=\sqrt{\alpha\beta}x.$
This leads to a rescaling of velocities as
$v^{\prime}$ $:=\sqrt{\frac{\beta}{\alpha}}v,$
$u^{\prime}$ $:=\sqrt{\frac{\beta}{\alpha}}u,$
and of the particle density as
$\rho^{\prime}:=\frac{\rho}{\sqrt{\alpha\beta}}.$
The rescaled noise term reads
$\eta^{\prime}:=\frac{\eta}{\sqrt{\alpha}}.$
Using these normalised variables, we obtain dimensionless Langevin equations of the form
[…]
$-\delta\sqrt{\frac{\beta^{3}}{\alpha}}\partial_{x^{\prime}}\rho^{\prime}+\sqrt{\frac{2D\beta}{\alpha^{2}}}\eta^{\prime}.$ (S.3)
D Minimisation of the autocorrelation time of collisions increases the discounted reward
[…]
First, remember that, we can express the expected discounted reward, $\langle G\rangle$, as a sum over the expected discounted reward conditioned on the reward observed at zero-th step, $R_{0}$,
$\langle G\rangle=\sum_{r\in\{r_{-},r_{+}\}}P(R_{0}=r)\langle G|R_{0}=r\rangle.$
Suppose that the particles are randomly distributed across the system, then the probability of $R_{0}$ being $r_{-}$, namely the probability of collision, is given by the average particle density, $\langle\rho\rangle$, hence
$\langle G\rangle=\langle\rho\rangle\langle G|R_{0}=r_{-}\rangle+(1-\langle\rho\rangle)\langle G|R_{0}=r_{+}\rangle.$ (S.4)
Notice that, by the definition of $G$, the conditional expectation, $\langle G|R_{0}=r\rangle$, is,
[…]
where $\tau$ is the autocorrelation time of the reward of type $r$. Combining Eq. (S.4) and Eq.  (S.5), and evaluating the integral, we obtain
$\langle G|R_{0}=r\rangle$ $=\frac{1}{\log(\gamma)}\left[\frac{\langle\rho\rangle\left(r_{-}\left(\langle\rho\rangle-\tau_{-}\log(\gamma)\right)-\langle\rho\rangle+1\right)}{\tau_{-}\log(\gamma)-1}-\frac{(\langle\rho\rangle-1)\left(\langle\rho\rangle r_{-}-\tau_{+}\log(\gamma)-\langle\rho\rangle+1\right)}{\tau_{+}\log(\gamma)-1}\right]$
$\sim C-\tau_{-}\langle\rho\rangle\left(1-\langle\rho\rangle\right)\left(1-r_{-}\right)+\mathcal{O}(\tau_{-}^{2})\qquad\text{as}\quad\tau_{-}\to 0$ (S.6)
where we set $r_{+}$ to 1 in equation (S.6), and $C$ is independent of the autocorrelation time of collisions $\tau_{-}$. As a result, decreasing $\tau_{-}$ will increase $\langle G\rangle$.

### v2

@@
 Current artificial intelligence systems show near-human-level capabilities when deployed in isolation.
-Systems of a few collaborating intelligent agents are being engineered to perform tasks collectively.
-This raises the question of whether robotic matter, where many learning and intelligent agents interact, shows emergence of collective behaviour.
-And if so, which kind of phenomena would such systems exhibit?
-Here, we study a paradigmatic model for robotic matter: a stochastic many-particle system in which each particle is endowed with a deep neural network that predicts its transitions based on the particles’ environments.
-For a one-dimensional model, we show that robotic matter exhibits complex emergent phenomena, including transitions between long-lived learning regimes, the emergence of particle species, and frustration.
-We also find a density-dependent phase transition with signatures of criticality.
-Using active matter theory, we show that this phase transition is a consequence of self-organisation mediated by emergent inter-particle interactions.
-Our simple model captures key features of more complex forms of robotic systems.
-The recent advances in artificial intelligence have brought about systems with near-human-level capabilities in some specific tasks.
-These advances mostly rely on deep neural networks, which can be trained efficiently using large data sets.
-In robotics, artificial intelligence is used to give robots the capability to interpret and manipulate their environment [1].
-Recently, artificial intelligence systems are being combined to achieve better performance compared to individual systems [2, 3, 4, 5, 6].
-In the context of robotics, such collectives contain multiple robotic units that are programmable and interact with their environment.
+Systems consisting of collaborating intelligent agents are being engineered to perform tasks collectively.
+This raises the question of whether robotic matter, where many learning and intelligent agents interact, shows the emergence of collective behaviour.
+And if so, what kind of phenomena would such systems exhibit?
+Here, we study a paradigmatic model for robotic matter: a stochastic many-particle system in which each particle is endowed with a deep neural network that optimizes its transitions based on the particles’ environments.
+For a one-dimensional model, we show that robotic matter exhibits complex phenomena arising from emergent interactions, including transitions between long-lived learning regimes, the emergence of particle species, and frustration.
+We also find an abrupt, density-dependent change in the behaviour of particles.
+Using active matter theory, we show that this phenomenon is a reflection of a phase transition with signatures of criticality.
+Our model captures key phenomena observed in more complex forms of robotic systems and identifies robotic matter as a laboratory for non-equilibrium physics.
+Recent advances in artificial intelligence have led to systems with near-human-level capabilities in some specific tasks [1, 2, 3, 4, 5, 6, 7].
+These advances rely primarily on deep neural networks, which can be trained efficiently using large datasets.
+In robotics, artificial intelligence is used to give robots the ability to interpret and manipulate their environment [8].
+Recently, artificial intelligence systems have been combined to achieve better performance compared to individual systems [9, 10, 11, 12, 13].
+In the context of robotics, such collectives comprise multiple programmable robotic units that interact with their environment.
 Programmed interactions enable such robot collectives to create different formations [14, 15, 16] and react to global signals to manipulate objects [17].
@@
 (c) Numerical protocol of stochastic updating and learning.
-Robot collectives have been experimentally implemented across spatial scales and system sizes  [11, 12].
-Due to the limitations imposed by miniaturization, there is a trade-off between the number of robots in a collective and the capacity of artificial intelligence that can be implemented in individual units.
+Robot collectives have been experimentally implemented across spatial scales and system sizes [18, 19, 14, 20, 21].
+Due to the limitations imposed by miniaturisation, there is a trade-off between the number of robots in a collective and the capacity of artificial intelligence that can be implemented in individual units.
 Macroscopic robot collectives comprise, for example, small groups of interacting robots which are each controlled by a performant artificial intelligence [22].
@@
 At the micron scale, robots interact via physical or chemical interactions between particles that are either self-driven active particles or particles driven by external forces and controlled via global signals.
-Such systems have, for example, been implemented as magnetic micro-disks driven by time-varying magnetic fields.
-Due to their small scale, the capabilities of individual particles in these systems are limited and they are therefore usually controlled externally using global signals.
-Conceptually, one can distinguish between different levels on which learning and execution occurs.
-In decentralised learning and decentralised execution, units learn and behave independently of each other [14].
-In the case of centralised learning and decentralised execution, units are controlled by a single machine-learning algorithm, which has been employed in active matter systems to learn chemotactic behaviour [15, 16].
-Finally, in the case of centralised learning and decentralised execution, where a single neural network is used by all particles individually but trained using the trajectory of all particles, 2D and 3D active intelligent swimmers moving collectively to reduce their energy expenditure  [17].
-Empirical studies in robot collectives have highlighted the emergence of subgroups of robots that may fulfil specialized tasks [18].Such speciation phenomena include the emergence of implicit leaders that provide direction of guidance to other robots.
-Such experiments have also highlight the emergence of deadlock states, robots perform suboptimal behaviour [19, 20, 21].
-Here, we study a paradigmatic theoretical framework for robotic matter that predicts key phenomena of empirical robot collectives.
-Specifically, we study stochastic many-particle systems in which each particle individually is endowed with artificial intelligence.
-In these systems the stochastic state transitions are defined by deep neural networks associated with each particle.
-The predictions by distinct deep neural networks are based on a particle’s environment and they learn by analysing the particle’s trajectory over time.
+Such systems have, for example, been implemented as magnetic micro-disks driven by time-varying magnetic fields [23, 24, 25].
+Due to their small scale, the capabilities of individual particles in these systems are limited, and they are therefore usually controlled externally using global signals.
+Conceptually, one can distinguish between different modes with which learning and execution occur  [26, 27, 28].
+In the mode of decentralised learning and decentralised execution, units can learn and behave independently of each other.
+This mode, for example, has been studied in cooperative and competitive games  [29, 30].
+In the mode of centralised learning and decentralised execution, units are controlled by a single machine-learning algorithm.
+This mode has been used, for example, in active matter systems to study chemotactic behaviour [31, 32, 33, 34], as well as to teach intelligent swimmers to collectively reduce their energy expenditure [35].
+Different modes give rise to different phenomenology.
+Empirical studies in robot collectives have highlighted the emergence of subgroups of robots that may fulfil specialised tasks [36].
+Such speciation phenomena include the emergence of implicit leaders that guide other robots.
+Such experiments have also highlighted the emergence of deadlock states, in which robots perform suboptimal behaviour [37, 38, 39, 40].
+Here, we study a paradigmatic theoretical framework for robotic matter that exhibits key phenomena of empirical robot collectives.
+Specifically, we study stochastic many-particle systems in which each particle has a distinct artificial intelligence.
+In these systems, the stochastic state transitions are defined by deep neural networks associated with each particle.
+The optimization done by distinct deep neural networks are based on a particle’s environment, and they learn by analysing the particle’s trajectory over time.
+Particles, therefore, can learn arbitrarily complex interactions.
 We find that these systems exhibit complex spatio-temporal behaviour: Over time, increasingly complex interactions between particles evolve in a succession of temporal learning regimes.
 We observe the evolution of multiple particle species and frustration, where the system is locked long-term in a suboptimal state.
-We also find a density-dependent phase transition with signatures of criticality.
-Using active matter theory, we show that this phase transition exhibits signatures of criticality and is a consequence of self-organisation mediated by emergent interactions between particles.
+Using active matter theory, we show that the system exhibits a density-dependent phase transition, with signatures of criticality, and that this transition is a consequence of self-organisation mediated by emergent interactions between particles.
 Figure 2: Representative trajectories showing the positions of particles over time in distinct temporal regimes (rows).
 Each column corresponds to a simulation at the specified particle density.
-For visual purposes we rescaled positions and time in the following way: The x-axis is the lattice position divided by the total lattice size (from left to right 2e4, 2e3, 1e3, 5e2, 4e2, 334 sites, respectively).
 The y-axis denotes the time in episodes divided by the typical time it takes for a particle to cross the system once.
-The starting time for each row is indicated in grey and the final time is on the top-right corner of each panel.
-Line colors show the sign of the average relative velocity of that particle in the last 2e5 episodes of the simulations.
+The starting time for each row is indicated in grey, and the final time is in the top-right corner of each panel.
+Line colours show the sign of the average relative velocity of that particle in the last 2e5 episodes of the simulations.
 Trajectories of representative particles are highlighted in bold.
-For densities 0.4, 0.5, and 0.6 we have randomly sampled 100 particles to make individual lines distinguishable.
+For densities 0.4, 0.5, and 0.6, we have randomly sampled 100 particles to make individual lines distinguishable.
 Results
@@
 These transition rates follow from the probabilities of each particle performing individual state transitions $x_{i}\to x_{i}^{\prime}$.
-In our model, each particle is endowed with a unique deep neural network (DNN) which takes the system configuration, $\{x_{i}\}_{i=1,\ldots,N}$, as an input and computes the single-particle transition rate as an output.
+In our model, each particle is endowed with a distinct deep neural network (DNN) which takes the system configuration, $\{x_{i}\}_{i=1,\ldots,N}$, as an input and computes the single-particle transition rate as an output.
 Because the system configuration includes the states of other particles, the DNNs may encode effective interactions between particles.
@@
 1(a).
-Here, to investigate the range of possible phenomenologies such systems may exhibit, we study the arguably simplest possible implementation of this framework.
+Here, to investigate the range of possible phenomenologies such systems may exhibit, we study, arguably, the simplest possible implementation of this framework.
 In this implementation, the state $x_{i}$ of particles is defined by their position on a one-dimensional lattice of $L$ sites with periodic boundary conditions, Fig.
@@
 This fixes the unit of time.
-These transition rates are determined by the output layer of feed-forward DNNs which take the lattice environment of size $l$ centered around a given particle as an input.
-The DNN parameters associated with a given particle are dynamically updated using deep reinforcement learning on the particles’ past trajectory to maximise a discounted reward.
+These transition rates are determined by the output layer of feed-forward DNNs, which takes the lattice environment of size $l$ centred around a given particle as an input.
+The DNN parameters associated with a given particle are dynamically updated using deep reinforcement learning on the particle’s past trajectory to maximise a discounted reward.
 The reward function gives a large negative reward when a particle moves to an already-occupied lattice site, and a small positive reward otherwise (Appendix A).
@@
 Group-agent reinforcement learning (GARL) is used for cases in which each agent operates in a separate environment and only communicates with others to share knowledge [48, 49].
-Here, we used an actor-critic reinforcement learning scheme with temporal-difference learning scheme to update the DNN parameters by maximising the cumulative discounted reward calculated from the particle trajectories in each episode  [31] (see Methods).
-All figures shown here were obtained from simulations, where the number of particles, $N$, was set to 200200200, the local environment size, $l$, was 212121, the system size, $L$, was set depending on particle density, and the discount factor, $\gamma$, was 0.90.90.9.
-DNNs had a triangular architecture with 21 neurons in the input layers, and 10, 5, and 2 neurons in the successive layers, respectively.
+Here, we used an actor-critic reinforcement learning scheme with a temporal-difference learning scheme to update the DNN parameters by maximising the discounted reward calculated from the particle trajectories in each episode [28] (see Methods).
+All figures shown here were obtained from simulations, where the number of particles, $N$, was set to $200$, the local environment size, $l$, was $21$, the system size, $L$, was set depending on particle density, and the discount factor, $\gamma$, was $0.9$.
+DNNs had a triangular architecture with 21 neurons in the input layer, and $21$, $10$, 55, and $2$ neurons in the successive layers, respectively.
 We simulated up to $5\cdot 10^{6}$ episodes with learning rate $5\cdot 10^{-4}$.
-We also tested DNNs with up to five layers, but we observed no qualitative differences in the simulations.
+We also tested DNNs with $21$ neurons in each layer, but we observed no qualitative differences in the simulations.
 To simulate dilute systems without reducing the number of particles, we increased the number of lattice sites to achieve a given particle density while keeping the size of the environment, $l$, fixed.
-This was to ensure that the DNN sizes and the training dynamics at different particle densities are consistent.
+This was done to ensure that the DNN sizes and the training dynamics at different particle densities are consistent.
 This fixes the units of length to the lattice spacing.
-From the definition of the model, several questions must be asked that in this form do not arise in active matter systems: First, particles appear in other particles’ environment.
+From the definition of the model, several questions arise that in this form do not arise in active matter systems: First, particles appear in other particles’ environment.
 Does this lead to the emergence of effective interactions between particles?
@@
 Despite the simplicity of the model, we find that particles form complex spatio-temporal structures after a short transient time.
-These structures change qualitatively over time, indicating that the dynamics transits through different temporal regimes.
-The spatio-temporal structures also depend on the particle density in that particles at low densities tend to align their velocity in the same direction, while above a density threshold of 0.3 particles at high densities move in opposite directions.
+These structures change qualitatively over time, indicating that the dynamics transition through different temporal regimes.
+The spatio-temporal structures also depend on the particle density in that particles at low densities tend to align their velocity in the same direction, while above a density threshold of 0.3 particles, at high densities move in opposite directions.
 In the following, we will analyse these changes as a function of the density more rigorously.
@@
 Particles are grouped by their asymptotic velocity, which is calculated as an average over the final 2e5 episodes of the simulations.
-The dashed lines on the time axis denote in this and in all following figure two characteristic times: $t_{1}\approx 1e4$ episodes marks the initial formation of global alignment at low densities; $t_{2}\approx 3e4$ episodes marks the initial formation of two groups of particles moving in opposite direction at high densities.
+The inlay show the density $0.5$ between episodes $5e3$ and $6e5$.
 (b) Average absolute velocity of all particles for two representative densities (colours defined in (c)).
+The dashed lines on the time axis denote in this and in all other figures two characteristic times: $t_{1}\approx 1e4$ episodes mark the initial formation of global alignment at low densities; $t_{2}\approx 3e4$ episodes mark the initial formation of two groups of particles moving in opposite directions at high densities.
 (c) Log2-fold change of the standard deviation of the number of right-moving particles across simulations.
-The lower horizontal dashed line is the variance expected if the direction of movement was assigned by chance.
+The lower horizontal dashed line is the variance expected if the direction of movement were assigned by chance.
 The upper dashed line is the maximal variance expected for globally aligned motion.
@@
 To this end, we first consider the relative velocity of particles with respect to each other.
-It is defined by the difference between the individual, instantaneous velocity of a given particle, $v_{i}$, and the average velocity in the system at a given time, $\expectationvalue{v_{i}}$, $v_{i}-\expectationvalue{v_{i}}$.
+It is defined by the difference between the individual, instantaneous velocity of a given particle, $v_{i}$, and the average velocity in the system at a given time, $\expectationvalue{v_{i}}$.
 We calculated the relative velocity for each particle and averaged over particles with the same sign of the asymptotic velocity.
 Figure 3(a) shows that, for low densities, the particles have vanishing relative velocities and therefore are moving in the same direction with a similar absolute velocity.
-Figure 3(b) shows that this occurs at a characteristic time $t_{1}\approx 1e4$ episods.
+Figure 3(b) shows that this occurs at a characteristic time $t_{1}\approx 1e4$ episodes.
 For densities larger than 0.3, particles form two groups moving in opposite directions.
 This could reflect a random assignment of the direction of movement of particles.
-However, the variance of particles moving in a given direction across all simulations is smaller than what is expected from a statistically independent assignment and predicted by the binomial for millions of episodes ($p=0.003$, F-test, Fig.
-3(c)).
-Significantly, this implies that, in the high-density regime, the direction of particles is not chosen randomly, but tuned into a precisely balanced proportion of left- and right-moving particles on the system level.
-The fine-tuning of particle directions to a 50-50 split at high densities seems counterintuitive given that they are rewarded for not colliding with other particles.
-Indeed, while the average reward at low concentrations increases monotonically and reaches an optimal value, at high densities the reward first increases and then decreases at a time that is comparable to the formation of the two groups of particles, Fig.
+However, the variance of particles moving in a given direction across all simulations is smaller than what is expected from a statistically independent assignment of direction ($p=0.003$, F-test, Fig.
+3(c)) and lasts for millions of episodes.
+Significantly, this implies that, in the high-density regime, the direction of particles is not randomly chosen, but tuned into a precisely balanced proportion of left- and right-moving particles on the system scale.
+The fine-tuning of particle directions to a balanced 50-50 split at high densities seems counterintuitive, given that they are punished for colliding with other particles.
+Indeed, while the average reward at low concentrations increases monotonically and reaches almost an optimal value, at high densities the reward first increases and then decreases at a time that is comparable to the formation of the two groups of particles, Fig.
 3(d).
-This observation raises two questions: first, what is the reason for the abrupt change in the behaviour of particles?
+This observation raises two questions: first, what is the reason for the abrupt change in the behaviour of particles as the density increases?
 And secondly, why is the high-density regime fine-tuned to a suboptimal reward?
-A potential answer to these questions could either be that particles learn different behaviours at high and low densities.
+A potential answer to these questions could be that particles learn different behaviours at high and low densities.
 This should then be reflected in the dynamics of the parameters of the DNNs.
@@
 In both density regimes, the nearest-neighbour distance decreases initially with time, showing that particles tend to cluster with other particles.
-The time evolution of the average nearest-neighbour distance is highly non-monotonic, reflecting different temporal regimes.
+However, the time evolution of the average nearest-neighbour distance is highly non-monotonic, reflecting different temporal regimes.
 Despite the formation of clusters, particles nevertheless reduce the probability of overlapping with other particles, Fig.
 4(b).
-To further understand the time evolution of spatial structures we calculated the pair correlation function, which gives the average density of particles as a function of the distance to a given particle, Fig.
+To further understand the time evolution of spatial structures, we calculated the pair correlation function, which gives the average density of particles as a function of the distance to a given particle, Fig.
 4(c).
@@
 Analysis of interactions
-To understand how the dynamics of the DNN parameters relate to the formation of spatio-temporal structures, we analyse the forces the DNNs encode.
-Intuitively, interactions between particles are encoded in the temporal change in its velocity, averaged over all environments.
-To develop this idea, we first represent a given environment by a vector $\vec{e}$, which contains the number of particles at each position in the environment.
-We now quantify the change in the velocity due to changes in the particle’s environment.
-To this end, we note that the change in the velocity can be written as the product of changes due to changing environments, and the temporal evolution of the distribution of finding a given environment.
-Interactions are encoded in the former part, which is given by the directional gradient of the velocity with respect to changes in the environment.
-In our case, the velocity is determined by the probability of hopping right in a given environment via $2p(\vec{e})-1$.
-Therefore, the directional derivative of the DNN output for a given input is a natural choice for quantifying interactions.
-To implement this, for a given environment $\vec{e}_{0}$, we quantified the change in the transition rates $p_{i}$ when the environment $\vec{e}_{0}$ is changed slightly by $\Delta e$.
-Here, we perturb the environment $\Delta\vec{e}$ by the addition of a single particle.
-With this, we can quantify interactions by calculating the difference in the neural network output between the original environment and the perturbed environment,
-$\frac{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}{\left|\Delta\vec{e}\right|}\,.$ (1)
-To quantify the type of interactions more qualitatively, we multiply (1) with the sign of the distance of the particle in $\Delta e$ to obtain a quantity that quantifies the degree with which particles are repulsed by the addition of a new particle to their environment (repulsiveness), $R=-\operatorname{sign}{(\Delta x)}\left[{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}\right]\big{/}{\left|\Delta\vec{e}\right|}$.
+To understand interactions between particles, for a given environment $\vec{e}_{0}$, we quantified the change in the transition probabilities $p_{i}$ when the environment $\vec{e}_{0}$ changes slightly by $\Delta e$.
+Here, we perturb the environment $\vec{e}_{0}$ by adding a single particle.
+With this, we can quantify interactions by calculating the difference in the neural network output between the original environment and the perturbed environment (see Appendix G),
+$\nabla_{\Delta\vec{e}}p(\vec{e_{0}})\approx\frac{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}{\left|\Delta\vec{e}\right|}.$ (1)
+To quantify the type of interactions more qualitatively, we multiply (1) with the sign of the distance of the particle in $\Delta e$ to obtain a quantity that quantifies the degree with which particles are repulsed by the addition of a new particle to their environment, which we term repulsiveness, $R=-\operatorname{sign}{(\Delta x)}\left[{p_{i}(\vec{e}_{0}+\Delta\vec{e})-p_{i}(\vec{e}_{0})}\right]\big/{\left|\Delta\vec{e}\right|}$.
 Negative values of the repulsiveness indicate attractive interactions.
@@
 As a first step towards quantifying interactions between particles, we therefore used a nonlinear dimensionality-reduction technique, uniform manifold approximation, and projection (UMAP), to obtain a first insight into the time evolution of interactions.
-Particles that are close in these projections tend to behave similarly in interactions.
+Particles that cluster these projections tend to behave similarly in interactions.
 The UMAPs show the systematic time evolution of the response of particles to changes in their environments, Fig.
 6(b).
-This time evolution is qualitatively different in the high and low density regimes, meaning that particles develop different interactions in both regimes, Fig.
+This time evolution is qualitatively different in the high and low-density regimes, meaning that particles develop different interactions in both regimes, Fig.
 6(b).
@@
 6(c), showing that particles consistently develop similar possible interactions across runs.
-At high densities, we observe the emergence of four clusters of particles in the projected space after a time $2\cdot 10^{5}$ which corresponds to the formation of the two groups of particles moving in different directions, Fig.
+At high densities, we observe the emergence of four clusters of particles in the projected space after a time $2\cdot 10^{5}$, which corresponds to the formation of the two groups of particles moving in different directions, as shown in Fig.
 2(c).
-This means that particles evolve into four distinct species, which are distinguished by the way they interact with their environments.
-Two pairs of these clusters identify particles moving ballistically in a given direction, Fig.
+This means that particles evolve into four distinct species that are distinguished by the way they interact with their environments.
+Two pairs of these clusters identify particles that move ballistically in a given direction, Fig.
 6(c).
 Particles moving in the same direction further separate into two groups, which interact in different ways depending on higher-order properties of the environments.
-We were not able to identify simple rules that distinguish between interactions of particles in the clusters corresponding to particles moving in the same direction, showing that particles develop complex interactions that do not follow patterns as in other systems studied in statistical physics.
+We were unable to identify simple rules that distinguish between interactions of particles in the clusters corresponding to particles moving in the same direction, showing that particles develop complex interactions that do not follow patterns as in other systems studied in statistical physics.
+The emergence of groups of particles that interact differently resembles speciation phenomena in evolutionary biology.
 To gain deeper insights into these interactions, we parametrised environments by the distance of the particle in $e_{j}$, $\Delta x$, and the distance of the first and second nearest neighbours in $\vec{e}_{0}$, $k_{1}$ and $k_{2}$, respectively.
@@
 This is expected by the choice of the reward function.
-Additionally, interactions tend be become longer in range over time, as particles obtain non-zero repulsion with respect to other particles at increasing distances.
-While particles at high densities only interact with the nearest and next-nearest neighbour sites particles at low densities interact over longer distances at large times.
-Figure 7(a) also shows characteristic branching points, that mark time points after which the response of particles in a given set of environments becomes non-homogeneous.
+Additionally, interactions tend to become longer in range over time, as particles obtain non-zero repulsion with respect to other particles at increasing distances.
+While particles at high densities only interact with the nearest and next-nearest neighbour sites, particles at low densities interact over longer distances at large times.
+Figure 7(a) also shows characteristic branching points that mark time points after which the response of particles in a given set of environments becomes non-homogeneous.
 At these time points, particles learn to distinguish an increasing number of different environments.
-To investigate how interactions depend on the base environment $\vec{e}_{0}$ we further parametrised the base environment by the distance to the first nearest neighbour, $k_{1}$, and the distance to the second nearest neighbor, $k_{2}$.
-At low densities, particles first distinguish between base environments, where $k_{1}=1$ and $k_{2}$ is either equal or larger than 2, Figure 7(b).
+To investigate how interactions depend on the base environment $\vec{e}_{0}$ we further parametrised the base environment by the distance to the first nearest neighbour, $k_{1}$, and the distance to the second nearest neighbour, $k_{2}$.
+At low densities, particles first distinguish between base environments, where $k_{1}=1$ and $k_{2}$ is either equal to or larger than 2, Fig.
+7(b).
 For low densities, after the emergence of nearest-neighbour repulsion, particles first learn to respond to the presence of a particle at a neighbouring site ($\Delta x=1$) depending on the position of the second nearest-neighbour particle.
 Specifically, particles learn to distinguish between environments, where the second nearest neighbour is two sites away compared to a longer distance.
-Later, particles learn to respond to the presence of a particle two sites away ($\Delta x=2$) depending on the position of a second nearest neighbour at the same distance.
+Later, particles learn to respond to the presence of a particle two sites away ($\Delta x=2$), depending on the position of a second nearest neighbour at the same distance.
 For high densities, after the emergence of nearest-neighbour repulsion, we observe a similar sequence of branching points but at later time points during the simulation.
 A large number of environments tend to follow similar trends, indicating that the DNNs in particles do not distinguish between these environments in terms of repulsiveness and that DNNs initially do not interpret the full complexity of the environment.
-At characteristic times that partially overlap with the boundaries of the temporal regimes defined on the macroscopic scale we observe branching points after which environments become distinguishable by the DNNs to increasing degrees.
+At characteristic times that partially overlap with the boundaries of the temporal regimes defined on the macroscopic scale, we observe branching points after which environments become distinguishable by the DNNs to increasing degrees.
 Therefore, over time, DNNs of particles learn to distinguish between an increasing number of environments at characteristic time points.
@@
 Each line shows the repulsiveness in a different environment.
-(a) Repulsiveness colored by the distance to the particle added to the base environments.
+The arrow and the dashed circles indicate the characteristic branching points of the interactions.
+(a) Repulsiveness colored by the distance to the particle was added to the base environments.
 (b) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments.
 Environments are coloured by the distance to the second nearest neighbour.
-(c) Repulsiveness for the case that a particle is added at distance $\Delta x=1$ to the base environments.
+(c) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments.
 Environments are coloured by the distance to the second nearest neighbour.
 Phenomenological theory
-With this, we are now in a position to come back to the question of what is the origin of the abrupt change in the behaviour of particle as a function of the particle density and the associated decrease of rewards.
-Our analyses have shown both: macroscopic order in space and time, which indicates collective behaviour of particles; and highly complex, density-dependent interactions that depend on evolving particle “species”.
-To test the hypothesis that the abrupt change in the behaviour of particles is a consequence of self-organization, we now develop a minimal, phenomenological model that only comprises a minimal set of interactions without explicit density dependence.
-To this end, we use the framework of active matter theory, and Langevin equations.
+We next address the origin of the abrupt change in the behaviour of the particle as a function of the particle density and the associated decrease in rewards?
+Our analyses have shown both macroscopic order in space and time, which indicates collective behaviour of particles, and highly complex, density-dependent interactions that depend on evolving particle “species”.
+To test the hypothesis that the abrupt change in the behaviour of particles is a consequence of self-organisation, we now develop a minimal, phenomenological model that comprises only a minimal set of interactions without explicit density dependence.
+To this end, we use the framework of active matter theory and Langevin equations.
 We consider the position $x_{i}(t)$, and velocity $v_{i}(t)$ of a particle $i$ at a given instance of time, $t$.
@@
 where $\gamma(v)$ is a generalized, non-equilibrium friction, $K_{i}$ is the force acting between particle $i$ and all other particles in its environment, $D$ is the noise amplitude, and $\eta$ is Gaussian white noise.
-We first identify the shape of the interactions $K_{i}$ from our simulations.
-To this end, we take a coarse-grained perspective and quantify the force acting on a particle $i$ for a given average velocity field $u(x)$ and density field $\rho(x)$ averaged in a small environment around the position $x_{i}$, $K_{i}(\rho,u)$.
-In our simulations, interactions evolve over time and are highly complex.
-They cannot be easily translated into simple interaction rules.
-For the purpose of our argument, we identify two density-, particle, and species-independent components of interactions: First, we have shown in Fig.
-7(b) that interactions between particles are predominantly repulsive.
-This gives rise to a contribution to the interaction term of the form $-\partialderivative*{\rho(x)}{x}\evaluated{}_{x=x_{i}}$.
+[…]
 Further, interactions aim to minimise not only the rate of collisions, but also their correlations.
@@
 S1).
-The autocorrelation time is proportional to the time a pair of particles spend in each other’s vicinity, such that particles minimise a function that is proportional to $1/|v-u|$.
+The autocorrelation time is proportional to the time a pair of particles spends in each other’s vicinity, such that particles minimise a function that is proportional to 1/|vu|1/|v-u|.
 This then leads to a force that is given by the derivative of this term with respect to $v$, $\operatorname{sign}(v-u)/|v-u|^{2}$.
@@
 Taken together, in our phenomenological description, this gives an effective force term for particle $i$ of the form
-$K_{i}=\zeta\rho\frac{\operatorname{sign}(v-u)}{|v-u|^{2}}-\delta\partialderivative{\rho}{x}\evaluated{}_{x=x_{i}}\,.$ (5)
-Figure 8: a) Representative trajectories showing the positions of particles at the specified particle density in the last 200 episodes.
-The colors show the sign of average velocity of the particles in the last 1e3 time points: red for positive and blue for a negative average velocity.
-b) Phase portrait of Eqs.
+[…]
+Figure 8: a) Representative trajectories from Langevin simulations of the equation (7) and (8), showing the positions of particles at the specified particle density in the last 3000 time points of the simulations, see Appendix B for details.
+The colours show the sign of the average velocity of particles in the last 100 time points (red: positive, blue: negative).
+For densities larger than 0.2, we have randomly sampled 256 particles to make individual lines distinguishable b) Phase portrait of Eqs.
 (9) and (10) for density 0.1.
 c) Phase portrait of Eqs.
-(9) and (10) for density 0.5.
-The colors show the sign of the velocity of the test particle in the Hartree approximation.
+(9) and (10) for a density 0.5.
+The colours show the sign of the velocity of the test particle in the Hartree approximation.
 d) Velocity autocorrelation time obtained from simulations of Eqs.
 (7) and (8).
-The dashed line shows the density 0.3.
 e) Velocity autocorrelation time obtained from simulation of the full model in Fig.
-1(b).
+1(b), see Appendix F for details.
+The dashed line shows the density of 0.3
 We now define the active friction term $\gamma(v)$.
 Because our system is out of equilibrium, the fluctuation-dissipation theorem does not hold, such that friction is not coupled to the noise amplitude.
-Instead, this term comprises a part that describes directed motion and a part that describes friction.
+Instead, this term comprises two parts: one that describes directed motion and another that describes friction.
 A phenomenological term that describes both is the Rayleigh-Helmholtz friction [50], which has been used to describe a variety of active matter systems.
@@
 Taken together, in non-dimensional form, the time-evolution of velocity and position of a given particle reads
-$\dot{x}=$ $v,,$ (7)
+$\dot{x}=$ $v\,$, (7)
 $\dot{v}=$ $v-v^{3}+s\frac{\rho\operatorname{sign}{(v-u)}}{(v-u)^{2}}-r\partialderivative{\rho}{x}+\sqrt{B}\eta,$ (8)
 where we have dropped the index $i$.
-We have non-dimensionalised the Langevin equations by rescaling time and space by $t^{\prime}=\alpha t$ and $x^{\prime}=x\sqrt{\alpha\beta}$, respectively, and removed the apostrophes to simplify the notation.
-The dimensionless parameters $s=\zeta\sqrt{\beta^{5}/\alpha^{3}}$ and $r=\delta\sqrt{\beta^{3}/\alpha}$ quantify the strength of interactions due to the minimisation of correlations and repulsion relative to the effect of self-propulsion, respectively.
+We have non-dimensionalised the Langevin equations by rescaling time and space by $t^{\prime}=\alpha t$ and $x^{\prime}=x\sqrt{\alpha\beta}$, respectively, and removed the apostrophes to simplify the notation (see Appendix C).
+The dimensionless parameters $s=\zeta\beta^{2}/\alpha^{2}$ and $r=\delta\sqrt{\beta^{3}/\alpha}$ quantify the strength of interactions due to the minimisation of correlations and repulsion relative to the effect of self-propulsion, respectively.
 $B=D\beta/\alpha^{2}$ is the dimensionless noise strength.
-In this model, the relative strength of the two interaction terms implicitly depends on the particle density.
-For high densities, fluctuations in the particle density are weak, such that we expect the term that minimises correlations to dominate.
+In this model, for high densities, fluctuations in the particle density are weak, such that we expect the term that minimises correlations to dominate.
 Intuitively, this leads to a maximisation of the difference between the particle’s velocity $v$ and the average velocity $u$, which could globally lead to a split as observed in our simulations.
-For low densities, the term that minimises correlations is negligible and we expect particle repulsion to be dominant.
+For low densities, the term that minimises correlations is negligible, and we expect particle repulsion to be dominant.
 To test this numerically, we performed stochastic simulations of Eqs.
 (7) and (8) using a finite difference method (see Appendix C).
-We indeed found that particles show global alignment at low densities and a split into groups moving in opposite direction for high densities (Fig.
+We indeed found that particles show global alignment at low densities and a split into groups moving in opposite directions for high densities (Fig.
 8).
-Again, the variance of particles moving in the positive direction across simulations is smaller than expected by chance ($p=2.22e^{-16}$), suggesting a collective tuning of the sizes of both groups.
-These observations suggest that the abrupt change in the behavior of particles in Fig.
-3(a) is a phase transition resulting from collective behavior induced by interactions between particles.
+Again, the variance of particles moving in the positive direction across simulations is smaller than expected by chance (t-test, $p=6.16e^{-7}$), suggesting a collective tuning of the sizes of both groups.
+These observations suggest that the abrupt change in the behaviour of particles, as seen in Fig.
+3(a), is a phase transition resulting from collective behaviour resulting from the effective interactions between particles.
 To investigate this more quantitatively, we analytically derived the coevolution of the velocity of a particle, $v$, and the average velocity of all particles $\bar{u}$.
-To this end, followed Ref.
-[33] and used a moment expansion up to second order in the average velocity to obtain a set of nonlinear differential equations,
-$\dot{v}=$ $(1-v^{2})v+s\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+r(\bar{u}-v),,$ (9)
+To this end, following Ref.
+[51], we used a moment expansion up to second order in the average velocity to obtain a set of nonlinear differential equations (see Appendix E),
+$\dot{v}=$ $(1-v^{2})v+s\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+r(\bar{u}-v)\,$, (9)
 $\dot{\bar{u}}=$ $\bar{u}-\bar{u}(\bar{u}^{2}+3T)\,.$ (10)
 Here, $T=\langle(v-\bar{u})^{2}\rangle$ is the variance of the velocities, which we estimated from our simulations of the Langevin equations.
-The co-evolution of the velocity of individual particles and the average velocity is summarized in phase portraits (Fig.
+The co-evolution of the velocity of individual particles and the average velocity is summarised in phase portraits (Fig.
 8).
@@
 (7) and (8).
-For low densities, we find two attractors with $v=\bar{u}$ that correspond to states, where the velocity of particles aligns with the average velocity in the system.
+For low densities, we find two attractors with $v=\bar{u}$ that correspond to states where the velocity of particles aligns with the average velocity in the system.
 For high densities, these attractors are located horizontally on a line given by $\bar{u}=0$.
@@
 This corresponds to the observed state with precisely tuned equal proportions of particles moving in either direction.
-Taken together, our analysis shows that the abrupt change of the behaviour as a function of the particle density is is an emergent phenomenon resulting from interactions between particles.
+Taken together, our analysis suggests that the abrupt change in the behaviour as a function of the particle density is an emergent phenomenon resulting from the effective interactions between particles.
 It is therefore a phase transition as in statistical physics systems.
-For high densities, interactions between particles have the macroscopic effect .
-In this regime, particles optimise their reward locally, without being able to predict the suboptimal consequences of their behaviour on the macroscopic scale.
-This resembles a learning analogon to frustration phenomena in disordered systems or the fallacy of the commons in social sciences.
+At all densities, interactions between particles have a macroscopic effect.
+In the high-density regime, particles optimise their reward locally, without being able to predict the suboptimal consequences of their behaviour on the macroscopic scale.
+This resembles a learning analogue to frustration phenomena in disordered systems or the fallacy of the commons in the social sciences.
 This analysis raises the question of whether this phase transition resembles known phase transitions in (non-equilibrium) statistical physics.
-To investigate this, we calculated autocorrelation functions of the velocities of individual particles using the batch means method [34], both in the Langevin simulations and in the simulations of the full model comprising intelligent agents, for different particle densities.
-In both cases, the autocorrelation time diverges at the density at which the phase transition occurs.
+To investigate this, we calculated autocorrelation functions of the velocities of individual particles, both in the Langevin simulations and in the simulations of the full model comprising intelligent agents, for different particle densities.
+In both cases, the autocorrelation time diverges at the density at which the phase transition occurs (Fig.
+8) .
 Diverging correlations are characteristic of critical points, suggesting critical behaviour.
 A further characterisation of the phase transition requires the computation of critical exponents, which is currently not feasible for intelligent agents.
-It is, however, of note that the phase transition occurs at a density, at which the range covered by the repulsive interactions of each particle multiplied by number of particles equals the lattice size, that is, the effective volume fraction is equal to 1.
+It is, however, of note that the phase transition occurs at a density at which the range covered by the repulsive interactions of each particle multiplied by the number of particles equals the lattice size; that is, the effective volume fraction is equal to 1.
 This could indicate a relation to jamming transitions.
@@
 Taken together, we studied a paradigmatic model of robotic matter.
-In this model, particles in a stochastic lattice gas are endowed with deep neural networks that take the spatial environment of a particle as an input and predict transition rates.
+In this model, particles in a stochastic lattice gas are endowed with deep neural networks that take the spatial environment of a particle as input and predict transition rates.
 Our extensive simulations reveal a rich phenomenology, including complex emergent phenomena leading to a phase transition, the evolution of multiple particle species, temporal learning regimes, and a phenomenon akin to frustration, where particles adopt suboptimal strategies.
 The phase transition we observe shows signs of criticality.
-Although it is not possible to determine critical exponents, the behaviour close to the phase transition point shows similarities to a jamming transition.
-Stochastic many-particle systems as the one studied in this work differ qualitatively from systems typically studied in active matter theory: because each particle has many more intrinsic degrees of freedom than the number of particles one can typically simulate on computers, they effectively lack microscopic symmetries.
-While in statistical physics one is typically interested in the breaking of microscopic symmetries, here the question arises if symmetries emerge as a consequence of learning, and whether these symmetries are then broken again.
-Because deep neural networks of sufficient size can, in principle, approximate arbitrarily complex function interactions between particles do not necessarily follow simple patterns.
+Although it is not possible to determine critical exponents, the behaviour close to the phase-transition point shows similarities to a jamming transition.
+Stochastic many-particle systems, such as the one studied in this work, differ qualitatively from systems typically studied in active matter theory: Because each particle has many more intrinsic degrees of freedom than the number of particles that one can typically simulate on computers, they effectively lack microscopic symmetries.
+While in statistical physics one is typically interested in the breaking of microscopic symmetries, here the question arises whether symmetries emerge as a consequence of learning, and whether these symmetries are then broken again.
+Because deep neural networks of sufficient size can, in principle, approximate arbitrarily complex function interactions between particles, they do not necessarily follow simple patterns.
 They may be qualitatively different between particles and environments.
-This raises the question of in how far physical theories are adequate in describing the time evolution of such systems.
+This raises the question of how far physical theories are adequate in describing the time evolution of such systems.
 In our work, particles indeed developed complex interactions that cannot be described by simple mathematical functions.
-However, a simple phenomenological theory focussing on rudimentary aspects of these interactions is able to predict a key aspect of the system, namely the existence of a density-dependent phase transition.
+However, a simple phenomenological theory focusing on the rudimentary aspects of these interactions can predict a key aspect of the system, namely the existence of a density-dependent phase transition.
+Reinforcement learning has been successfully used to control the behaviour of non-equilibrium systems, by controlling the motion of a single particle subject to hydrodynamic forces [20, 52], by exerting global control on many-particle systems [53, 54], or by enforcing alignment by rewarding high particle densities [29].
+In our work, the use of neural networks allows for the learning of almost arbitrarily complex relations between the environment of a particle and its behaviour.
+These possibilities are explored and used by the particles, as evidenced in the learning of increasingly complex interactions and the formation of particle species.
+This again leads to remarkably rich spatio-temporal behaviour.
 Although our model is a drastic simplification of real robotic systems, our results show similarities to much more complex robotic systems.
 For example, Ref.
-[19] studied a hide and seek game between intelligent agents, in which agents passed through different, metastable learning regimes and interactions between agents lead to a decrease in the collected reward.
-Apart from interacting robots, also virtual artificial intelligence systems, like the widely-used large-language systems, are now beginning to be combined in order to achieve better performance [35].
+[37] studied a hide-and-seek game between intelligent agents, in which agents passed through different, metastable learning regimes, and interactions between agents led to a decrease in the collected reward.
+In addition to interacting robots, virtual artificial intelligence systems, such as widely used large language systems, are now beginning to be combined to achieve better performance [55].
 This raises the question of how to design these systems to achieve an expected outcome.
-In some cases, interacting learning objectives straightforwardly translate into collective behaviour [14].
-Our work implies that learning can be drastically influenced in unpredictable manners by collective behaviour on the macroscopic scale.
+In some cases, learning objectives straightforwardly translate into collective behaviour [29].
+Our work implies that learning can be drastically influenced in unpredictable ways by collective behaviour on the macroscopic scale.
 Acknowledgements
@@
 A.A., O.B., and S.R.
-conceptualized the work.
+conceptualised the work.
 A.A.
@@
 supervised the work; O.B.
-wrote the simulation code and analysed O.B.
+wrote the simulation code and analysed data.
+O.B.
 performed analytical calculations.
@@
 wrote the initial draft of the manuscript.
-All authors edited and approved of the final version of the manuscript.
+All authors edited and approved the final version of the manuscript.
+DATA AVAILABILITY
+The data are available from the authors upon reasonable request.
 A Implementation of the stochastic simulations
 […]
-$\sum_{i=1}^{n_{j}-1}-A_{i}\log{p_{i}},$
-where $n_{j}$ is the number of transitions the particle $j$ made during the episode, and $A_{i}:=R_{i}+\gamma V_{i+1}-V_{i}$, and $p_{i},R_{i},V_{i}$ are the probability with which the particle made the transition, the reward it got and the value predicted by the critic at step $i$, respectively.
-To train the critic, we used the loss $\sum_{i=1}^{n_{j}-1}A_{i}^{2}/(n_{j}-1)$.
-Figure S1: The average autocorrelation time of collisions over 101010 replicates.
-The shaded area around the curves denote the standard error.
-The averages are calculated at logarithmically samples time points.
+[…]
+Where $n_{j}$ is the number of transitions the particle $j$ made during the episode, and $\delta_{i}:=R_{i}+\gamma V_{i+1}-V_{i}$, and $p_{i},R_{i},V_{i}$ are the probability with which the particle made the transition, the reward it got and the value predicted by the critic at step $i$, respectively, where $i$ run from $1$ until $(n_{j}-1)$.
+To train the critic, we used the loss $\sum_{i=1}^{n_{j}-1}\delta_{i}^{2}/(n_{j}-1)$.
+Figure S1: The rescaled average autocorrelation time of collisions, averaged over $10$ replicates.
+The autocorrelation time for each density is rescaled by their value at episode 11.
+The shaded area around the curves denotes the standard error.
+The averages are calculated at logarithmically spaced time points.
 B Numerical solution of the Langevin equations
-We used the Euler scheme to integrate the Langevin equations.
-We used 200200200 particles with time increment $\delta t=0.01$, and the system size was determined based on $L=200/\langle\rho\rangle$ where $\langle\rho\rangle$ is the average particle density.
+We used the Euler–Maruyama scheme to integrate the non-dimensionalised Langevin equation.
+We tested $N=200$ and $N=1024$ particles with time increment $\delta t=0.001/\tau_{t}$, where $\tau_{t}=1/\alpha$ is the characteristic time-scale, and for a total time $T=1\times 10^{5}/\tau_{t}$.
+The system size was determined based on $L=N/\langle\rho\rangle$ where $\langle\rho\rangle=\rho_{0}/\tau_{l}$ is the average non-dimensional particle density and $\tau_{l}=1/\sqrt{\alpha\beta}$ is the characteristic length-scale.
 We used periodic boundary conditions.
-We set $s=1.17,r=12.50,B=0.02$, and we used a local radius of 555 to calculate the local density field $\rho(x)$ and the local velocity field $u(x)$ for each particle.
-We run 151515 to 242424 replicates per particle density.
-To calculate the gradients of the local fields, we subtracted value of the field in the right of the particle at distance 555 from the value of the field in the left of the particle at distance 555, and divided this by 555.
+We set $\alpha=16,\beta=1,\delta=50,\zeta=75,D=2.5$, and we set the radius of interactions, $R$, to $1/\tau_{l}$ to calculate the local density and velocity field $\rho(x),u(x)$.
+For $N=200$, we run $15$ to $24$ replicates per particle density for $N=200$, while for $N=1024$, we run 2 replicates per particle density due to the limited availability of GPU hardware.
+To prevent divergences in the numerical integration of the singular term $\frac{\operatorname{sign}{(v^{\prime}-u^{\prime})}}{(v^{\prime}-u^{\prime})^{2}}$, we regularised it with the $\tanh$ function.
 C Non-dimensionalisation of the Langevin equation
-To non-dimensionalize the Langevin equations, we rescaled time and positions as
+To non-dimensionalise the Langevin equations, we rescaled time and positions as
 $t^{\prime}$ $:=\alpha t$
@@
 D Minimisation of the autocorrelation time of collisions increases the discounted reward
-[…]
-First, remember that, we can express the expected discounted reward, $\langle G\rangle$, as a sum over the expected discounted reward conditioned on the reward observed at zero-th step, $R_{0}$,
+In this appendix, we show that the expected discounted reward, $\langle G\rangle$, increases as the autocorrelation time of the collisions $\tau_{-}=\int_{0}^{\infty}\langle\delta(R_{0}-r_{-})\cdot\delta(R_{t}-r_{-})\rangle\differential t d$ecreases.
+First, remember that we can express the expected discounted reward, $\langle G\rangle$, as a sum over the expected discounted reward conditioned on the reward observed at the zero-th step, $R_{0}$,
 $\langle G\rangle=\sum_{r\in\{r_{-},r_{+}\}}P(R_{0}=r)\langle G|R_{0}=r\rangle.$
@@
 Notice that, by the definition of $G$, the conditional expectation, $\langle G|R_{0}=r\rangle$, is,
-[…]
-[…]
+$\langle G|R_{0}=r\rangle$ $=\int_{0}^{\infty}dt\gamma^{t}\langle R_{t}|R_{0}=r\rangle$
+$=\int_{0}^{\infty}dt\gamma^{t}\left(\langle R_{t}\rangle+e^{-t/\tau}(r-\langle R_{0}\rangle)\right),$ (S.5)
 where $\tau$ is the autocorrelation time of the reward of type $r$. Combining Eq. (S.4) and Eq.  (S.5), and evaluating the integral, we obtain
-$\langle G|R_{0}=r\rangle$ $=\frac{1}{\log(\gamma)}\left[\frac{\langle\rho\rangle\left(r_{-}\left(\langle\rho\rangle-\tau_{-}\log(\gamma)\right)-\langle\rho\rangle+1\right)}{\tau_{-}\log(\gamma)-1}-\frac{(\langle\rho\rangle-1)\left(\langle\rho\rangle r_{-}-\tau_{+}\log(\gamma)-\langle\rho\rangle+1\right)}{\tau_{+}\log(\gamma)-1}\right]$
+$\langle G\rangle$ $=\frac{1}{\log(\gamma)}\left[\frac{\langle\rho\rangle\left(r_{-}\left(\langle\rho\rangle-\tau_{-}\log(\gamma)\right)-\langle\rho\rangle+1\right)}{\tau_{-}\log(\gamma)-1}-\frac{(\langle\rho\rangle-1)\left(\langle\rho\rangle r_{-}-\tau_{+}\log(\gamma)-\langle\rho\rangle+1\right)}{\tau_{+}\log(\gamma)-1}\right]$
 $\sim C-\tau_{-}\langle\rho\rangle\left(1-\langle\rho\rangle\right)\left(1-r_{-}\right)+\mathcal{O}(\tau_{-}^{2})\qquad\text{as}\quad\tau_{-}\to 0$ (S.6)
 where we set $r_{+}$ to 1 in equation (S.6), and $C$ is independent of the autocorrelation time of collisions $\tau_{-}$. As a result, decreasing $\tau_{-}$ will increase $\langle G\rangle$.
+E Coevolution of single-particle velocity $v$ and the average velocity $\bar{u}$
+We follow the approach taken by references [51, 56] and use the moment-expansion method to find a closed set of equations for the moments of the probability density $P(x,v,t)$ of finding a particle at position $x$ with velocity $v$ at time $t$.
+We then truncate this hierarchy of equations at order two to arrive at an approximation of the true probability density.
+First, to be able to obtain a closed set of equations, we start by replacing the term $r\partial_{x}\rho$ in equation (8) with the term r(uv)r(u-v).
+This replacement allows us to express $\dot{v}$ in terms only of $v$ and $u$.
+Next, we define our moment-expansion hierarchy.
+Recall that, n-th moment, $M_{n}(x,t)$,
+$M_{n}(x,t):=\left\langle v^{n}\right\rangle=\frac{1}{\rho}\int v^{n}P(x,v,t)dv,\quad n>0$ (S.7)
+The normalisation $\rho$ is the zeroth moment, which is equivalent to the marginal density
+$M_{0}(x,t)=\rho(x,t)=\int P(x,v,t)dv$ (S.8)
+Multiplying the nth moment by the density and taking the derivative with respect to time, we arrive at the dynamics of the nth moments of the velocity
+$\frac{\partial}{\partial t}\left(M_{0}M_{n}\right)=\int v^{n}\frac{\partial P}{\partial t}dv.$ (S.9)
+Next, we use an effective single-particle description and use the Fokker-Planck equation to obtain
+$\frac{\partial P}{\partial t}=$ $-v\frac{\partial}{\partial x}P+B\frac{\partial^{2}}{\partial v^{2}}P$ (S.10)
+$-\frac{\partial}{\partial v}\left\{\left(1-v^{2}\right)v+s\frac{\rho\operatorname{sign}{(v-u)}}{(v-u)^{2}}+r\left(u-v\right)\right\}P$
+Then Inserting equation (S.10) into equation (S.9), and using $\lim_{v\to\pm\infty}P(x,v,t)=0$, the terms containing partial derivatives with respect to $v$ can be partially integrated, resulting in
+$\frac{\partial}{\partial t}\left(M_{0}M_{n}\right)=$ $-\frac{\partial}{\partial x}\rho\left\langle v^{n+1}\right\rangle+n(n-1)B\rho\left\langle v^{n-2}\right\rangle$ (S.11)
+$+n\rho\left(\left\langle v^{n}\right\rangle-\left\langle v^{n+2}\right\rangle\right)$
+$+n\rho\left(s\rho\int v^{n-1}\frac{v-u}{|v-u|^{3}}Pdv\right)$
+$+n\rho\left(r\left\langle v^{n-1}\right\rangle-\left\langle v^{n}\right\rangle\right).$
+We now express $v(x,t)=\bar{u}(x,t)+\delta v(x,t)$, as the average velocity plus deviations from it. We assume that $\langle\delta v^{l}\rangle=0$ for odd exponents $l=1,3,5,$... . Thus we obtain
+$\langle v\rangle$ $=\bar{u}$ (S.12)
+$\langle v^{2}\rangle$ $=\bar{u}^{2}+T$ (S.13)
+$\langle\delta v^{2}\rangle$ $=T.$ (S.14)
+The dynamics of the first moment are given by
+$\dot{\bar{u}}=\bar{u}-\bar{u}(\bar{u}^{2}+3T)$ (S.15)
+In the limit where the number of particles goes to infinity, one can approximate $u(x,t)$ with $\bar{u}(x,t)$.
+Furthermore, thanks to the translational invariance of the system, we can treat the fields $v(x,t),\bar{u}(x,t)$ as scalar variables $v(t),u(t)$.
+Now, embedding a tracer particle whose dynamics are given by equation (8) into a system whose dynamics are determined by equation (S.15), we obtain the following set of two equations.
+$\dot{\bar{u}}$ $=\bar{u}-\bar{u}(\bar{u}^{2}+3T)$ (S.16)
+$\dot{v}$ $=v-v^{3}+\zeta\frac{\beta^{2}}{\alpha^{2}}\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+\delta\sqrt{\frac{1}{\alpha}}(\bar{u}-v)$ (S.17)
+which qualitatively capture the phenomenology of the system with varying density.
+Ideally, one can truncate the hierarchy at order three and above to obtain the dynamics of $T$.
+However, it was not possible to express the integral $\int v^{n-1}\frac{\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}Pdv$ in terms of the moments of $v$ for $n>1$ to close to hierarchy.
+As a result, we measured the value of $T$ at different densities from the agent-based simulations and used those values in Fig.
+8.
+F Calculation of the velocity autocorrelation function
+To calculate the velocity autocorrelation of individual particles in the full model comprising intelligent agents (Fig.
+8(e)), we used the batch means method [57].
+To calculate the velocity autocorrelation of individual particles in the many-body Langevin equation, we calculated the connected correlation function
+$=\langle v_{i}(t+\Delta t)-v_{i}(t)\rangle-\langle v_{i}(t)\rangle\langle v_{i}(t+\Delta t)\rangle$ (S.18)
+for each particle $i$. We then calculated the autocorrelation time using its definition
+$\tau_{i}=\int_{0}^{T}d(\Delta t)C_{i}(\Delta t)/C_{i}(0).$ (S.19)
+where $T$ was set to $500$.
+G Gradient of the neural networks as a proxy for interactions
+To understand how the dynamics of the DNN parameters relate to the formation of spatio-temporal structures, we analyse the forces the DNNs encode.
+Intuitively, interactions between particles are encoded in the temporal change in their velocity, averaged over all environments.
+To develop this idea, we first represent a given environment by a vector $\vec{e}$, which contains the number of particles at each position in the environment.
+To quantify the change in the velocity due to changes in the particle’s environment, we note that the change in the velocity, namely the force, can be written as
+$F$ $=\langle\dot{v}\rangle$
+[…]
+Using the conservation of probability, we can replace $\partial_{t}P(\vec{e},t)$ with the divergence of the probability flux, $j$, giving us
+$F$ $=-\int_{[0,\infty)^{l}}d\vec{e}(2p(\vec{e})-1)\nabla\cdot\vec{j}.$
+Now, using integration by parts, we obtain
+[…]
+The first term in the integral can be turned into a surface integral via the divergence theorem, giving us
+[…]
+$=0,$ (S.22)
+since, on the boundaries of the domain $0,\infty)^{l}$, the term $((2p(\vec{e})-1)\vec{j})\cdot\hat{n}$ vanishes either because $\vec{j}\cdot\hat{n}$ vanishes (in the case when all except one of $e_{i}$ is zero) or $(2p(\vec{e})-1$ goes to zero as $|\vec{e}|\to\infty$.
+As a result, finally we obtain
+[…]
+This motivates the use $\nabla p(\vec{e})$ as a proxy for the interactions of the particles.